OpenStudy (anonymous):

Please help will fan and medal!! what is the equation in standard form of a parabola that contains the following points (-2 -20) , (0,4) , (4, -20)

OpenStudy (anonymous):

y=-3x^2+6x+4

OpenStudy (anonymous):

how did you get that? @tosin

OpenStudy (anonymous):

ax² + bx + c = y gives you 3 equations: a(-2)² + b(-2) + c = -20 ∙ ∙ → 4a - 2b + c = -20 a(0)² + b(0) + c = 4 ∙ ∙ ∙ ∙ → c = 4 a(4)² + b(4) + c = -20 ∙ ∙ ∙ → 16a + 4b + c = -20 use eq2 to simplify eq1 and eq3 and get 4a - 2b = 16 ∙ ∙ ∙ → 2a - b = -12 16a + 4b = 40 ∙ → 4a + b = -6 add and get 6a = 18 ∙ ∙ ∙ a = 3